Matthew is 34 years older than Steven. 10 years ago, Matthew was three times Stevens age then. How old are they now
Answer by stanbon(75887)

Matthew is 34 years older than Steven. 10 years ago, Matthew was three times Stevens age then. How old are they now
---------
Equations:
m = s + 34
m-10 = 3(s-10)
-----
Substitute for "m" and solve for "s":
s+34 - 10 = 3s - 30
s+24 = 3s - 30
2s = 54
s = 27 yrs (Stephen's age now)
------
Solve for "m"
m = s + 34
m = 27+34
m = 61 yrs (Matthew's age now)